( 81/36) -3/4 x (( 25/9 ) -3/2 / ( 5/2) -3 )

( 81/36)-3/4 x  (( 25/9 )-3/2   /   ( 5/2) -3 )

Grade:9

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
We have:
\left (\frac{81}{36} \right )^{-\frac{3}{4}}\times \left ( \frac{\left ( \frac{25}{9} \right )^{-\frac{3}{2}}}{\left ( \frac{5}{2} \right )^{-3}} \right )=\left (\frac{36}{81} \right )^{\frac{3}{4}}\times \left ( \frac{\left ( \frac{9}{25} \right )^{\frac{3}{2}}}{\left ( \frac{2}{5} \right )^{3}} \right )=\left (\frac{\left (\sqrt{6} \right )^3}{3^3} \right )\times\left ( \frac{\left ( \frac{3}{5} \right )^3}{\left ( \frac{2}{5} \right )^{3}} \right )=\left (\frac{\left (\sqrt{6} \right )^3}{2^3} \right )=\frac{3}{2}\sqrt{\frac{3}{2}}Regards
Sumit